The owner of a candy store wishes to mix candy selling for $1.50 per pound with nuts selling for $1.00 pere pound to obtain a party mix to be sold for $1.20 per pound. How many pounds of each must be used to obtain 50 pounds of the mixture?
Answer by jorel1380(3719)

50x1.20=60
m+n=50
1.5m+1n=60
1.5m+1.5n=75
.5n=15
n=15/.5=30
m=20
The owner needs 20 lbs at $1.50/lb. and 30 lbs. at $1.00/lb to get the desired mixture..